OpenStudy (anonymous):

4 log u - 5 log u to the power of 6

OpenStudy (anonymous):

What is it you want to do with this?

OpenStudy (anonymous):

Write as a single logarithm

OpenStudy (anonymous):

ok. Use 2 principles about logs. 1) x[log y] = log (y^x) and 2) log x - log y = log (x/y).

OpenStudy (anonymous):

=log u ^4 -log u^5 =log(u^4 / u^5) =log(1 / u)